Power:

Stan and Olga are considering buying a sheep farm. A surveyor has supplied them with the given accurate sketch. Find the area of the property

Power:

 Not to scale

Moon:

yikes. not 100% sure on this but you could start by using law of cosines to find the missing sides, then using Heron's formula for the areas :S

Power:

What I tried to do: (Find area of left triangle) \(\large {A=\frac{1}{2}sq}SinP\) \(A=\frac12(8)(10)Sin(70º)\) \(A=45.10\) \(A=\frac12(s)(r)Sin(30º)\)

Power:

That second one I am stuck on

Power:

Oh

Power:

\(p^2=s^2+q^2-2sqCos(P)\) \(p^2=8^2+12^2-2(8)(12)Cos(70)\) \(p^2=142.33\) (Approximately) \(p=11.93\)?

Power:

Oh wow I got it now, thanks!